# Sparse autoencoders reveal selective remapping of visual concepts during adaptation

### v1

@@ 1 Introduction
Foundation models excel at fast adaptation to new tasks and domains with limited extra training data [27, 4, 34]. In the space of vision-language models, CLIP [27] became an important backbone for numerous applications [20, 18, 30]. The CLIP model consists of two transformer networks to encode text and image inputs. Various parameter-efficient adaptation techniques, such as adopting learnable tokens, have been proposed targeting either of the systems. While early works targeted the text encoder [38, 37], more recently it was shown that joint adaptation of both the text and image encoders can further improve classification performance [15, 16]. Despite these advances in adaptation methods, it remains an open question of how representations in a foundation model change during adaptation.
Recently, Sparse Autoencoders (SAEs; 3, 6) have gained emerging attention as a tool of mechanistic interpretability [33, 36] after showing its effectiveness in widely used LLMs [3, 19]. SAEs map dense model representations, which are difficult to interpret because multiple unrelated concepts are entangled (polysemantic), to sparse and interpretable (monosemantic) concepts.
@@
We use our PatchSAE to shed light on the internal mechanisms of foundation models during adaptation tasks. Recent state-of-the-art adaptation methods for CLIP [37, 38, 15, 16] add trainable, dataset specific tokens for adaptation, akin to a system prompt in LLMs. Through extensive analysis, we reveal a wide range of interpretable concepts of CLIP, including simple visual patterns to high-level semantics, employing our PatchSAE as an interpretability tool (Fig. 1(b)). We also localize the recognized concepts through token-wise inspections of SAE latent activations, while extending it to image-, class-, and task-wise understandings. Furthermore, we demonstrate that the SAE latents have a crucial impact on the model prediction in classification tasks through ablation studies. Lastly, we show evidence that prompt-based adaptation gains the performance improvement by tailoring the mapping between recognized concepts and the learned task classes (Fig. 1(c)).
@@ 2 Related Work
Mechanistic interpretability [11] aims to interpret how neural networks infer their outputs. To achieve this, it is natural to seek a deeper understanding of which feature (concept) is recognized by each neuron in the neural network [25]. For instance, the logit lens [24] approach attempts to understand the intermediate layer output by mapping it into the final classification or decoding layer. However, understanding neurons in human interpretable form is challenging due to the polysemantic [10] nature of neurons, where each neuron activates for multiple unrelated concepts. This property is attributed to superposition [10], where neural networks represent more features than the number of dimensions.
To overcome the superposition phenomenon in neural network interpretation, sparse autoencoders (SAEs) [31, 3] have recently gained significant attention. SAEs decompose model activations into a sparse latent space, representing them as dictionary of high dimensional vectors. Several studies [36, 6] have applied SAEs to language models. Using SAEs, Templeton [33] discovered bias- and safety-related features in large language models (LLMs), demonstrating that these features can be steered to alter the behavior of LLMs. Recent research extended the application of SAEs to vision-language models, such as CLIP [27]. Fry [13] and Daujotas [7] extracted interpretable concepts from the vision encoder of CLIP and Daujotas [8] utilized these features to edit image generation in a diffusion model. Rao et al. [29] named the SAE concepts using word embeddings from the CLIP text encoder and used them for a concept bottleneck model.
@@
Adapting vision-language foundation models like CLIP through fine-tuning requires large datasets and significant computation resources. In addition, the generalization ability of the model may be compromised after fine-tuning. As an alternative, prompt-based adaptation has recently emerged, training only a few learnable tokens while keeping the weight of pre-trained models fixed. CoOp [38] proposed adapting CLIP in few-shot learning setting by optimizing learnable tokens in the language branch, while Bahng et al. [1] applied prompt adaptation to the vision branch. MaPLe [15] improved few-shot adaptation performance by jointly adding learnable tokens to both the vision and language branches and considered as a base structure for more recent multimodal prompt adaptation methods [16]. Although these studies demonstrate that prompt learning is effective for adapting CLIP, there is still a lack of research focusing on how and why prompt learning enables such adaptation.
Our work focuses on exploring the CLIP image encoder using an SAE on a vision transformer. We choose MaPLe [15] as a representative structure of multimodal prompt adaptation and investigate the internal work of adaptation methods.
@@ 3.1 PatchSAE Architecture and Training Objectives
SAEs typically consist of a single linear layer encoder, followed by a ReLU non-linear activation function, and a single linear layer decoder [3, 6]. To train an SAE on CLIP Vision Transformer (ViT), we hook intermediate layer outputs from the pre-trained CLIP ViT and use them as self-supervised training data. We leverage all tokens including class (CLS) and image tokens from the residual stream output  [^1] The residual stream is the sum of the attention block’s output and its input, see Fig. 9(c). (Fig. 9(c)) of an attention block and feed them to the SAE. Formally, we take ViT hook layer output as an SAE input $\mathbf{z}$, multiply it with the encoder layer weight $W_{E}\in\mathbb{R}^{d_{\text{ViT}}\times d_{\text{SAE}}}$, pass to the ReLU activation $\phi$, then multiply with the decoder layer weight $W_{D}\in\mathbb{R}^{d_{\text{SAE}}\times d_{\text{ViT}}}$ [^2] We use bias terms for linear layers and centralize $\mathbf{z}$, i.e., $\text{SAE}(\mathbf{z}-\mathbf{b}_{\text{dec}}).$. The column (or row) vectors of the encoder (or decoder), $d_{\text{SAE}}$ vectors size of $\mathbb{R}^{d_{\text{ViT}}}$, correspond to the candidate concepts, i.e., SAE latent directions. We call the output vector of the activation layer (size of $\mathbb{R}^{d_{\text{SAE}}}$) as SAE latent activations. For simplicity, we use $f$ for the encoder and $g$ for the decoder:
@@ 3.2 Analysis method and evaluation setup
Summary statistics of SAE latents. To inform the general trend of an SAE latent, we compute summary statistics of the activation distribution [3]. We use the activated frequency and the mean activation values over a subset of training images. Using the class label information from a classification benchmark dataset such as ImageNet, we compute the label entropy [13] and standard deviation from the reference images. Specifically, we compute and interpret the statistics as follows:
@@
  • Label standard deviation. In ImageNet, class labels are organized in a hierarchical structure based on WordNet’s semantic relationships [9, 22]. We leverage this label structure and use the label standard deviation of reference images as a clue for the semantic granularity besides the label entropy when exploring the latents. We discuss more in §A.1.
@@
Localizing patch-level SAE latent activations. PatchSAE allows localizing an active latent in the image space. We treat the patch-level latent activation as a soft segmentation mask. Precisely, given an image $\mathbf{x}_{i}$ and an SAE latent index $s$, we multiply each patch $\mathbf{x}_{i,j}$ with the corresponding latent activation value $\mathbf{h}_{i,j}[s]$ for visualization. For example in Fig. 2(c), we highlight “yellow flowers” or “dogs with black, white, and brown colors” concepts from the input image. Separating the patches relevant to the targeting latent from the input and reference images shows a clearer view of the concept [^3] We recommend trying on/off segmentation mask option in the interactive demo.
@@ 4.1 Impact of SAE Latents on Classification
We explore the influence of SAE latents on the final model prediction in classification tasks. We replace the intermediate layer representation of CLIP image encoder with the SAE reconstructed output to steer the model output [^4] We add reconstruction score when replacing the intermediate layer output with SAE reconstructed one following [33]. [33] by selectively using a subset of SAE concepts. Then we compute cosine similarity between text and image encoder outputs for the classification task. Fig. 5(a) summarizes the procedure.
@@ 4.1.1 Analysis method and experiment setup
Zero-shot classification. We conduct ImageNet-1K [9] zero-shot classification using OpenAI CLIP ViT-B/16 (§Reproducibility) with an ensemble of 80 OpenAI ImageNet templates [27] to compute text features for each class and conduct classification by computing cosine similarity between image features out text features (Fig. 5(a)).
@@ 4.1.2 Key findings
The results for the top-$k$ SAE latent masking experiments are shown in Fig. 5(b)&(c). Using all SAE latents (on all; identity mask) recovers the original classification performance (i.e., using the original model representation without replacing it with the SAE output) with small reconstruction errors (64.82% for the identity mask and 68.25% for the original). Using the all-zero mask (off all) removes all relevant information, and hence results in the accuracy of 0.1%. Ablating randomly selected or task-wise latents do not show significant affect to the classification accuracy until we use sufficient number of latents. On the other hand, ablating the per-class top activating latents shows a crucial impact on classification performance. We observe a clear performance improvement or degradation in accordance with the increased or decreased number of active SAE latents, respectively. The results of this analysis show that some SAE latents contain rich information that is critical for class discrimination. Moreover, the search for such latents can be narrowed down to the top activating SAE latents that are frequently activated across inputs with the same ground-truth class.
@@ 4.2.1 Analysis method and experiment setup
Base-to-novel classification. Following the setup introduced by [38], we split the downstream task dataset classes into two groups and consider the first half as base and the remaining as novel classes, then conduct classification on two groups separately. We use total 11 benchmark datasets (§Reproducibility): ImageNet-1K [9], Caltech101 [12], OxfordPets [26], StanfordCars [17], Flowers102 [23], Food101 [2], FGVC Aircraft [21], SUN397 [35], DTD [5], EuroSAT [14], and UCF101 [32].
Prompt-based adaptation methods append learnable tokens to a frozen pretrained CLIP and train the added tokens on downstream tasks. Specifically, MaPLe [15] adds learnable tokens at the input layer and the first few layers both for text and image encoders (Fig. 5(a)). In the base-to-novel setting, MaPLe uses few-shot samples from each of the base classes to train the learnable tokens. We use officially released MaPLe weights (§Reproducibility) for the experiments.
@@
Figure 6: Different impact of SAE latents. We show two example cases of Flowers102 (left) and EuroSAT (right). In each figure, we show two example images of the ground-truth (GT) class, a scatter plot view of class-level SAE latent activations comparison ($x$-axis: CLIP, $y$-axis: MaPLe trained on the base classes of the downstream task), reference images from ImageNet and the downstream dataset for top latents, and the top-$k$ masking results as a prediction heatmap. We show reference images with segmentation mask in the EuroSAT case.
@@ 6 Conclusion
We used publicly available model checkpoints for CLIP (link) and MaPLe (link). We used OpenAI ImageNet templates for zero-shot classification (link). We will open source our code, SAE model weights and raw results for secondary analysis upon publication of the paper. We only used publicly available datasets following the official implementation of MaPLe (dataset descriptions) which are cited in the main text.
@@ Acknowledgments
HsL was supported through a DAAD/NRF fellowship in the NRF Summer Institute Programme, 2024 (57600422). This work was supported by the Helmholtz Association’s Initiative and Networking Fund on the HAICORE@KIT partition.
@@ A.1 Training performance
Quantitative metrics. We follow the literature on SAEs [3, 33] to set quantitative metrics. The mean squared error (MSE) loss indicates reconstruction ability. Contrastive loss with and without SAE informs the reconstruction ability as well. Close contrastive losses indicate that the SAE reconstructs the input better. The L1 loss and the L0 metric indicate the sparsity of SAE. The lower the values are, the less number of SAE latents are activated for the given input. We start by reproducing training performance as reported in previous studies [13]. Then we ablated training hyperparameters $\lambda_{l_{1}}$, expansion factor, ghost gradient technique [^5] Ghost gradient is introduced as an improvement of neuron resampling [3] that addresses dead neurons due to sparsity regularization of training., and the initialization of decoder bias term (mean vs. geometric median of the training dataset). Furthermore, we ablated model architectures (ViT-B/16 and ViT-L/14), training tokens (CLS vs. all; Fig. 8), and hook layers (§A.2).
@@
Supplements to summary statistics. In addition to the frequency and the mean value of activation distribution [3], we use the label entropy [13] and the label standard deviation that can give an intuition about concept granularity. The label standard deviation is tailored for a labeled dataset such as ImageNet, where the label structure contains a hierarchical structure of English words. In this case, the standard deviation indicates whether the latent is capturing a distinct label from ImageNet dataset or other attributes such as the style (or domain) or patterns of image. For example, the dog latent might be activated by different breeds of dogs, so the number of unique labels is high (high entropy) but the gap between labels might be low (low standard deviation). On the other hand, blue color latent might be activated by diverse blue objects or scenes whose labels can be very far away (high entropy and high standard deviation).
Although the quantitative metrics of reconstruction and sparsity validate that SAE is trained as intended, they do not provide rich information about the validity and interpretability of SAEs. Therefore, we utilize SAE latent summary statistics and reference images for qualitative evaluation. We mostly follow the configurations as selected by Fry [13], confirming that the chosen setup shows reasonable performance. To be compatible with both zero-shot and adapted method MaPLe, which releases official weight on ViT-B/16, we choose model architecture as CLIP ViT-B/16. For a deeper understanding, we use all image tokens in addition to the CLS token. We also note recent progress of SAE architectures and training techniques such as gated SAE [28], using SAE on other components’ output (such as attention output or MLP output), but we focus on the base architecture of SAEs [3] treating the advanced techniques as out-of-scope.
Table 2: SAE training configurations and performance. We share results for six design choices: ViT hooking layer, training tokens, ViT architecture, L1 coefficient $\lambda_{l_{1}}$, expansion factor ($d_{\text{ViT}}\times\text{(expansion factor)}=d_{\text{SAE}}$), using ghost gradient technique, and initialization of decoder bias term. Default configurations are colored as gray. CL indicates the contrastive loss. * indicates baseline result [13]. Italic indicates the ablated configuration.
@@
Figure 8: SAE statistics comparisons on training token ablation. Training SAE only for (a) the class token and for (b) all image tokens (including the class token) show a difference in statistics. Both have two clusters in the bottom left and top right regions but (b) showed a denser cluster in the top right region than (a). The plotting style is referenced from [13].

### v2

@@ 1 Introduction
Foundation models excel at fast adaptation to new tasks and domains with limited extra training data (Radford et al., 2021; Caron et al., 2021; Touvron et al., 2023). In the space of vision-language models, CLIP (Radford et al., 2021) became an important backbone for numerous applications (Liu et al., 2024; Li et al., 2023; Rombach et al., 2022). The CLIP model consists of two transformer networks to encode text and image inputs. Various parameter-efficient adaptation techniques, such as adopting learnable tokens, have been proposed targeting either of the systems. While early works targeted the text encoder (Zhou et al., 2022b, a), more recently it was shown that joint adaptation of both the text and image encoders can further improve classification performance (Khattak et al., 2023a, b). Despite these advances in adaptation methods, it remains an open question of how representations in a foundation model change during adaptation.
Recently, Sparse Autoencoders (SAEs; Bricken et al., 2023; Cunningham et al., 2023) have gained emerging attention as a tool of mechanistic interpretability (Templeton, 2024; Yun et al., 2021) after showing its effectiveness in widely used LLMs (Bricken et al., 2023; Lieberum et al., 2024). SAEs map dense model representations, which are difficult to interpret because multiple unrelated concepts are entangled (polysemantic), to sparse and interpretable (monosemantic) concepts.
@@
We use our PatchSAE to shed light on the internal mechanisms of foundation models during adaptation tasks. Recent state-of-the-art adaptation methods for CLIP (Zhou et al., 2022a, b; Khattak et al., 2023a, b) add trainable, dataset specific tokens for adaptation, akin to a system prompt in LLMs. Through extensive analysis, we reveal a wide range of interpretable concepts of CLIP, including simple visual patterns to high-level semantics, employing our PatchSAE as an interpretability tool (Fig. 1(b)). We also localize the recognized concepts through token-wise inspections of SAE latent activations, while extending it to image-, class-, and task-wise understandings. Furthermore, we demonstrate that the SAE latents have a crucial impact on the model prediction in classification tasks through ablation studies. Lastly, we show evidence that prompt-based adaptation gains the performance improvement by tailoring the mapping between recognized concepts and the learned task classes (Fig. 1(c)).
@@ 2 Related Work
Mechanistic interpretability (Elhage et al., 2021) aims to interpret how neural networks infer their outputs. To achieve this, it is natural to seek a deeper understanding of which feature (concept) is recognized by each neuron in the neural network (Olah et al., 2020). For instance, the logit lens (Nostalgebraist, 2020) approach attempts to understand the intermediate layer output by mapping it into the final classification or decoding layer. However, understanding neurons in human interpretable form is challenging due to the polysemantic (Elhage et al., 2022) nature of neurons, where each neuron activates for multiple unrelated concepts. This property is attributed to superposition (Elhage et al., 2022), where neural networks represent more features than the number of dimensions.
To overcome the superposition phenomenon in neural network interpretation, sparse autoencoders (SAEs) (Sharkey et al., 2022; Bricken et al., 2023) have recently gained significant attention. SAEs decompose model activations into a sparse latent space, representing them as dictionary of high dimensional vectors. Several studies (Yun et al., 2021; Cunningham et al., 2023) have applied SAEs to language models. Using SAEs, Templeton (2024) discovered bias- and safety-related features in large language models (LLMs), demonstrating that these features can be steered to alter the behavior of LLMs. Recent research extended the application of SAEs to vision-language models, such as CLIP (Radford et al., 2021). Fry (2024) and Daujotas (2024a) extracted interpretable concepts from the vision encoder of CLIP and Daujotas (2024b) utilized these features to edit image generation in a diffusion model. Rao et al. (2024) named the SAE concepts using word embeddings from the CLIP text encoder and used them for a concept bottleneck model.
@@
Adapting vision-language foundation models like CLIP through fine-tuning requires large datasets and significant computation resources. In addition, the generalization ability of the model may be compromised after fine-tuning. As an alternative, prompt-based adaptation has recently emerged, training only a few learnable tokens while keeping the weight of pre-trained models fixed. CoOp (Zhou et al., 2022b) proposed adapting CLIP in few-shot learning setting by optimizing learnable tokens in the language branch, while Bahng et al. (2022) applied prompt adaptation to the vision branch. MaPLe (Khattak et al., 2023a) improved few-shot adaptation performance by jointly adding learnable tokens to both the vision and language branches and considered as a base structure for more recent multimodal prompt adaptation methods (Khattak et al., 2023b). Although these studies demonstrate that prompt learning is effective for adapting CLIP, there is still a lack of research focusing on how and why prompt learning enables such adaptation.
Our work focuses on exploring the CLIP image encoder using an SAE on a vision transformer. We choose MaPLe (Khattak et al., 2023a) as a representative structure of multimodal prompt adaptation and investigate the internal work of adaptation methods.
@@ 3.1 PatchSAE Architecture and Training Objectives
SAEs typically consist of a single linear layer encoder, followed by a ReLU non-linear activation function, and a single linear layer decoder (Bricken et al., 2023; Cunningham et al., 2023). To train an SAE on CLIP Vision Transformer (ViT), we hook intermediate layer outputs from the pre-trained CLIP ViT and use them as self-supervised training data. We leverage all tokens including class (CLS) and image tokens from the residual stream output  [^1] The residual stream is the sum of the attention block’s output and its input, see Fig. 9(c). (Fig. 9(c)) of an attention block and feed them to the SAE. Formally, we take ViT hook layer output as an SAE input $\mathbf{z}$, multiply it with the encoder layer weight $W_{E}\in\mathbb{R}^{d_{\text{ViT}}\times d_{\text{SAE}}}$, pass to the ReLU activation $\phi$, then multiply with the decoder layer weight $W_{D}\in\mathbb{R}^{d_{\text{SAE}}\times d_{\text{ViT}}}$ [^2] We use bias terms for linear layers and centralize $\mathbf{z}$, i.e., $\text{SAE}(\mathbf{z}-\mathbf{b}_{\text{dec}}).$. The column (or row) vectors of the encoder (or decoder), $d_{\text{SAE}}$ vectors size of $\mathbb{R}^{d_{\text{ViT}}}$, correspond to the candidate concepts, i.e., SAE latent directions. We call the output vector of the activation layer (size of $\mathbb{R}^{d_{\text{SAE}}}$) as SAE latent activations. For simplicity, we use $f$ for the encoder and $g$ for the decoder:
@@ 3.2 Analysis method and evaluation setup
Summary statistics of SAE latents. To inform the general trend of an SAE latent, we compute summary statistics of the activation distribution (Bricken et al., 2023). We use the activated frequency and the mean activation values over a subset of training images. Using the class label information from a classification benchmark dataset such as ImageNet, we compute the label entropy (Fry, 2024) and standard deviation from the reference images. Specifically, we compute and interpret the statistics as follows:
@@
  • Label standard deviation. In ImageNet, class labels are organized in a hierarchical structure based on WordNet’s semantic relationships (Deng et al., 2009; Miller, 1995). We leverage this label structure and use the label standard deviation of reference images as a clue for the semantic granularity besides the label entropy when exploring the latents. We discuss more in §A.1.
@@
Localizing patch-level SAE latent activations. PatchSAE allows localizing an active latent in the image space. We treat the patch-level latent activation as a soft segmentation mask. Precisely, given an image $\mathbf{x}_{i}$ and an SAE latent index $s$, we multiply each patch $\mathbf{x}_{i,j}$ with the corresponding latent activation value $\mathbf{h}_{i,j}[s]$ for visualization. For example in Fig. 2(c), we highlight “yellow flowers” or “dogs with black, white, and brown colors” concepts from the input image. Separating the patches relevant to the targeting latent from the input and reference images shows a clearer view of the concept [^3] We recommend trying the on/off segmentation mask option in the interactive demo.
@@ 4.1 Impact of SAE Latents on Classification
We explore the influence of SAE latents on the final model prediction in classification tasks. We replace the intermediate layer representation of CLIP image encoder with the SAE reconstructed output to steer the model output [^4] We add reconstruction score when replacing the intermediate layer output with SAE reconstructed one following Templeton (2024). (Templeton, 2024) by selectively using a subset of SAE concepts. Then we compute cosine similarity between text and image encoder outputs for the classification task. Fig. 5(a) summarizes the procedure.
@@ 4.1.1 Analysis method and experiment setup
Zero-shot classification. We conduct ImageNet-1K (Deng et al., 2009) zero-shot classification using OpenAI CLIP ViT-B/16 (§Reproducibility) with an ensemble of 80 OpenAI ImageNet templates (Radford et al., 2021) to compute text features for each class and conduct classification by computing cosine similarity between image features out text features (Fig. 5(a)).
@@ 4.1.2 Key findings
The results for the top-$k$ SAE latent masking experiments are shown in Fig. 5(b,c). Using all SAE latents (on all; identity mask) recovers the original classification performance (i.e., using the original model representation without replacing it with the SAE output) with small reconstruction errors (64.82% for the identity mask and 68.25% for the original). Using the all-zero mask (off all) removes all relevant information, and hence results in the accuracy of 0.1%. Ablating randomly selected or task-wise latents do not show significant affect to the classification accuracy until we use sufficient number of latents. On the other hand, ablating the per-class top activating latents shows a crucial impact on classification performance. We observe a clear performance improvement or degradation in accordance with the increased or decreased number of active SAE latents, respectively. The results of this analysis show that some SAE latents contain rich information that is critical for class discrimination. Moreover, the search for such latents can be narrowed down to the top activating SAE latents that are frequently activated across inputs with the same ground-truth class.
@@ 4.2.1 Analysis method and experiment setup
Base-to-novel classification. Following the setup introduced by Zhou et al. (2022b), we split the downstream task dataset classes into two groups and consider the first half as base and the remaining as novel classes, then conduct classification on two groups separately. We use total 11 benchmark datasets (§Reproducibility): ImageNet-1K (Deng et al., 2009), Caltech101 (Fei-Fei et al., 2004), OxfordPets (Parkhi et al., 2012), StanfordCars (Krause et al., 2013), Flowers102 (Nilsback & Zisserman, 2008), Food101 (Bossard et al., 2014), FGVC Aircraft (Maji et al., 2013), SUN397 (Xiao et al., 2010), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), and UCF101 (Soomro, 2012).
Prompt-based adaptation methods append learnable tokens to a frozen pretrained CLIP and train the added tokens on downstream tasks. Specifically, MaPLe (Khattak et al., 2023a) adds learnable tokens at the input layer and the first few layers both for text and image encoders (Fig. 5(a)). In the base-to-novel setting, MaPLe uses few-shot samples from each of the base classes to train the learnable tokens. We use officially released MaPLe weights (§Reproducibility) for the experiments.
@@
Figure 6: Different impact of SAE latents. We show two example cases of Flowers102 (left) and EuroSAT (right). In each figure, we show two example images of the ground-truth (GT) class, a scatter plot view of class-level SAE latent activations comparison ($x$-axis: CLIP, $y$-axis: MaPLe trained on the base classes of the downstream task), reference images from ImageNet and the downstream dataset for top latents, and the top-$k$ masking results as a prediction heatmap. We show reference images with segmentation masks for EuroSAT.
@@ 6 Conclusion
We used publicly available model checkpoints for CLIP (link) and MaPLe (link). We used OpenAI ImageNet templates for zero-shot classification (link). Code, model weights and raw results are available at https://github.com/dynamical-inference/patchsae. We only used publicly available datasets following the official implementation of MaPLe (dataset descriptions) which are cited in the main text.
@@ Acknowledgments
HsL was supported through a DAAD/NRF fellowship in the NRF Summer Institute Programme, 2024 (57600422). This work was supported by the Helmholtz Association’s Initiative and Networking Fund on the HAICORE@KIT partition. This work was supported by Institute for Information & communications Technology Planning & Evaluation (IITP) and National Research Foundation of Korea (NRF) grants funded by the Korea government (MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST), No. 2022R1A5A7083908, and No. RS-2025-00555621).
@@ A.1 Training performance
Quantitative metrics. We follow the literature on SAEs (Bricken et al., 2023; Templeton, 2024) to set quantitative metrics. The mean squared error (MSE) loss indicates reconstruction ability. Contrastive loss with and without SAE informs the reconstruction ability as well. Close contrastive losses indicate that the SAE reconstructs the input better. The L1 loss and the L0 metric indicate the sparsity of SAE. The lower the values are, the less number of SAE latents are activated for the given input. We start by reproducing training performance as reported in previous studies (Fry, 2024). Then we ablated training hyperparameters $\lambda_{l_{1}}$, expansion factor, ghost gradient technique [^5] Ghost gradient is introduced as an improvement of neuron resampling Bricken et al. (2023) that addresses dead neurons due to sparsity regularization of training., and the initialization of decoder bias term (mean vs. geometric median of the training dataset). Furthermore, we ablated model architectures (ViT-B/16 and ViT-L/14), training tokens (CLS vs. all; Fig. 8), and hook layers (§A.2).
@@
Supplements to summary statistics. In addition to the frequency and the mean value of activation distribution (Bricken et al., 2023), we use the label entropy (Fry, 2024) and the label standard deviation that can give an intuition about concept granularity. The label standard deviation is tailored for a labeled dataset such as ImageNet, where the label structure contains a hierarchical structure of English words. In this case, the standard deviation indicates whether the latent is capturing a distinct label from ImageNet dataset or other attributes such as the style (or domain) or patterns of image. For example, the dog latent might be activated by different breeds of dogs, so the number of unique labels is high (high entropy) but the gap between labels might be low (low standard deviation). On the other hand, blue color latent might be activated by diverse blue objects or scenes whose labels can be very far away (high entropy and high standard deviation).
Although the quantitative metrics of reconstruction and sparsity validate that SAE is trained as intended, they do not provide rich information about the validity and interpretability of SAEs. Therefore, we utilize SAE latent summary statistics and reference images for qualitative evaluation. We mostly follow the configurations as selected by Fry (2024), confirming that the chosen setup shows reasonable performance. To be compatible with both zero-shot and adapted method MaPLe, which releases official weight on ViT-B/16, we choose model architecture as CLIP ViT-B/16. For a deeper understanding, we use all image tokens in addition to the CLS token. We also note recent progress of SAE architectures and training techniques such as gated SAE (Rajamanoharan et al., 2024), using SAE on other components’ output (such as attention output or MLP output), but we focus on the base architecture of SAEs (Bricken et al., 2023) treating the advanced techniques as out-of-scope.
Table 2: SAE training configurations and performance. We share results for six design choices: ViT hooking layer, training tokens, ViT architecture, L1 coefficient $\lambda_{l_{1}}$, expansion factor ($d_{\text{ViT}}\times\text{(expansion factor)}=d_{\text{SAE}}$), using ghost gradient technique, and initialization of decoder bias term. Default configurations are colored as gray. CL indicates the contrastive loss. * indicates baseline result Fry (2024). Italic indicates the ablated configuration.
@@
Figure 8: SAE statistics comparisons on training token ablation. Training SAE only for (a) the class token and for (b) all image tokens (including the class token) show a difference in statistics. Both have two clusters in the bottom left and top right regions but (b) showed a denser cluster in the top right region than (a). The plotting style is referenced from Fry (2024).